### v1

# DESI Dark Energy Time Evolution is Recovered by Cosmologically Coupled Black Holes

Abstract
Recent results from baryon acoustic oscillation measurements by the Dark Energy Spectroscopic Instrument (DESI) have found preliminary evidence that dark energy (DE) evolves with time, as parameterized by a $w_{0}w_{a}$ equation of state. In this study, we point out that the DESI-preferred $w_{0}w_{a}$ models are recovered when a $w_{0}w_{a}$ model is fit to DE produced by baryon conversion in cosmologically coupled black holes (BHs). This recovery does not require any ad hoc parameter adjustments; it solely depends on the independently measured cosmic star formation rate density. We discuss our result in the context of the missing baryon problem and the anomalously low sum of neutrino masses preferred by DESI. The global evolution of DE is an orthogonal probe of cosmological coupling, complementing constraints on BH mass growth from elliptical galaxies, stellar binaries, globular clusters, the LIGO-Virgo-KAGRA merging population, and X-ray binaries. A DE density that correlates with star formation is a natural outcome of cosmological coupling in BH populations.
1 Introduction
It has been nearly a quarter century since two pioneering experiments, the Supernova Cosmology Project [1] and the High-Z Supernovae Search Team [2], independently discovered the accelerated expansion of the universe using type Ia supernovae as standard candles. The acceleration was attributed to a pervasive form of energy, known as dark energy (DE) [3], whose nature has remained elusive.
In 2006, the Dark Energy Task Force (DETF) [4] proposed a comprehensive four-stage experimental program to investigate the nature of DE by determining whether the accelerated expansion is consistent with a cosmological constant, $\Lambda$, whether it evolves, or whether it arises from modifications to General Relativity (GR). For dynamic models of DE, the DETF adopted a simple linear parameterization of the equation of state of DE, known as the $w_{0}w_{a}$ parameterization:
$w=w_{0}+w_{a}(1-a)\,,$ (1.1)
where $a$ is the cosmological scale factor, and $w_{0}$ and $w_{a}$ are constants. To evaluate the stage of ongoing and proposed experiments, the DETF introduced the reciprocal of the area of the error ellipse enclosing the 95% confidence limit in the $w_{0},w_{a}$ plane as a Figure of Merit. The results from Stage III experiments [5, 6, 7] have shown consistency with a cosmological constant but exhibit small ($\lesssim 2.5\sigma$) deviations from $\Lambda$CDM in the $w_{0}>-1$, $w_{a}<0$ direction.
Recently, the Dark Energy Spectroscopic Instrument (DESI), the first Stage IV spectroscopic survey, reported significantly enhanced cosmology constraints from Baryon Acoustic Oscillation (BAO) measurements using its first year of data (DESI-1YR) [8]. The DESI collaboration finds that adopting a $w_{0}w_{a}$ DE equation of state and combining DESI-1YR BAO data with CMB lensing data and Type Ia SNe data from the Pantheon+, Union3 and DES-5YR datasets leads to deviations from $\Lambda$CDM by $2.5\sigma$, $3.5\sigma$ and $3.9\sigma$, respectively. The values of $w_{0}=-0.831\pm{0.066},-0.64\pm{0.011}$ and $-0.724\pm{0.071}$ and $w_{a}=-0.73^{+0.32}_{-0.28},-1.30^{+0.45}_{-0.39}$ and $-1.06^{+0.35}_{-0.31}$ for the combinations DESI + CMB + Pantheon+, Union3 and DES-5YR respectively, indicate a strong preference for $w_{0}>-1$ and large $w_{a}<0$. Substituting these values into the $w_{0}w_{a}$ parameterization, Eq. 1.1 implies that, at some point in the past, $w<-1$.
An equation of state $w<-1$ is often referred to as “phantom,” because it appears to generate energy densities from nothing. [^1] Phantom models typically require a negative kinetic energy term, which can be envisioned as a scalar field rolling up its potential [9]. In addition, phantom models have been argued to have theoretical inconsistencies [10, 11, 12, 13]. Because this would imply an unstable vacuum, the interpretation of DESI data within the context of the $\Lambda$CDM model [14, 15, 16, 17, 18, 19], as well as suggestions for interacting DE [20], quintessence scalar fields [21, 22, 23], and other scenarios beyond $\Lambda$CDM [24, 25], has attracted recent interest from the scientific community. Phantom DE is also considered a useful phenomenological description of models in which DE actually receives energy from another species. Models have been proposed where dark matter (DM) couples to DE [26, 27, 28, 29, 30, 31]. However, very little is known about DM, and it has yet to be detected directly [32, 33, 34, 35, 36], which currently limits the predictive power of these models for the phenomenology of DE.
Another possibility involves models where baryons couple to DE. For example, the conversion of DE into baryons explains the origin of the Big Bang in the inflationary paradigm [37, 38]. Conversely, the conversion of baryons into DE was proposed by Gliner to occur during the gravitational collapse of dead stars, as described in the paper that theoretically discovered what we now call DE [39]. Following this study, a diverse community of researchers has developed exact GR solutions of BHs with DE interiors [40, 41, 42, 43, 44] and explored condensed-matter and quantum field-theoretic interpretations (see e.g. [45, 46, 47]). These BH solutions often mimic the familiar Schwarzschild and Kerr BH models but are singularity-free and do not have horizons. This can lead to surprising phenomena when these objects are generalized to cosmological boundary conditions. For instance, there are known GR solutions for BHs [e.g. 48, 49] that expand in lockstep with their embedding cosmology, gaining mass independently of accretion or merger. [^2] Recent theoretical objections to cosmological coupling [50, 51, 52, 53, 54] demonstrate the existence of decoupled solutions. Because GR is a non-linear theory, the existence of a solution does not imply uniqueness, (see e.g. §1.2 of [55]) and so cannot exclude other solutions. Solutions with realistic cosmological boundary conditions and mass growth $\propto a$ establish these phenomena as a robust GR prediction [49, 56]. Coupling also follows from a mathematically rigorous treatment of the Einstein-Hilbert action for perturbed Robertson-Walker spacetime, where metric degrees of freedom are necessarily constrained in Fourier space [57, 58, 59].
Recently, observational evidence for purely cosmological mass growth has been reported in the super-massive black holes (SMBHs) of elliptical galaxies [60]. The preferred rate of mass growth for this population, $m\propto a^{3}$ [61] is consistent with observations from NANOgrav [62] and low-mass X-ray binaries (XRBs) [63]. Because cosmological number densities decrease $\propto 1/a^{3}$, if SMBHs gain mass $\propto a^{3}$, then (in the absence of production) their physical density is constant. Conservation of stress energy then requires that they enter Friedmann’s equations as DE, consistent with the DE-interior BH models. However, BH growth $\propto a^{3}$ exceeds the upper bounds $\propto a$, from the known coupled GR solutions and recent measurements in stellar-mass BH populations of LIGO-Virgo-KAGRA [64, 65, 66, 67], globular clusters [68], Gaia binaries [69], and high-mass XRBs [70]. Local measurements of cosmological mass growth are complicated by accretion and merger, and require accurate estimation of BH formation time, making orthogonal probes of this GR effect valuable for understanding the emerging theoretical and observational landscape.
DESI and forthcoming Stage IV surveys [71, 72, 73, 74, 75, 76, 77, 78] can provide positive evidence for DE within cosmologically coupled BHs. The cosmological background offers an averaged dynamics, free from selection effects and assumptions about minimum BH mass [^3] For BH models without horizons, the Tolman-Oppenheimer-Volkoff stability limit $\lesssim 2.2M_{\odot}$ for maximal neutron star mass can no longer be used to establish a lower-bound BH mass. This also confounds interpretation of low-mass remnants without optical or merger tidal-deformation signatures, e.g. [79]. required to constrain coupled mass growth in isolated stellar binaries. If BHs contribute as a DE species, then DE time-evolution will track BH production and growth. [^4] The possible role of stellar-collapse BHs in cosmic expansion was first proposed in the Gravitational Aether Theory, an extension to GR [80, 81]. Because stellar-collapse is expected to be the dominant BH formation channel over the DESI-sensitive redshift range $z\lesssim 3$, and this range overlaps with the peak rate of cosmic star formation around $z\sim 2$, DESI is well-positioned to search for a cosmologically coupled BH population. In this paper, we present the first search for this signal at “cosmic noon.” Because the unprocessed DESI-1YR data has not yet been publicly released, we analyze DE as sourced by cosmologically coupled stellar-collapse BHs (BH DE) within the $w_{0}w_{a}$ framework. We find that the $w_{0}w_{a}$ evolution expected from BH DE matches the DESI most-likely $w_{0}w_{a}$ models when those models include any SNe datasets.
Figure 1: Agreement between the $w_{0}w_{a}$ best-fit to the BH DE model and all combinations of DESI datasets. (Dotted) The $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a$ prediction of the BH DE model, where $w:=-1$ and DE is sourced by non-singular stellar collapse BHs, as described in Eq. 2.11. DESI uses a $w_{0}w_{a}$ parameterization, which limits its ability to fit predictions of the BH DE model. (Dashed) The $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a$ prediction given by the least-squares best-fit $w_{0}w_{a}$ parameterization to the BH DE model. (Solid) The $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a$ implied by DESI measurements of $w_{0}$ and $w_{a}$. Each panel is determined from the DESI-1YR central values of $w_{0}$, $w_{a}$, $\Omega_{m}$, and $H_{0}$ for DESI + CMB + Pantheon+ (top), DESI + CMB + Union3 (middle), and DESI + CMB + DESY5 (bottom) (see Table 3 in [82]). In all panels, the best-fit DESI model closely tracks the least-squares best-fit to the BH DE model with a standard star formation rate density [83]. Therefore, DESI observations are consistent with $w=-1$, provided that DE is produced by stellar-collapse to non-singular, cosmologically coupled, BHs.
2 Theory
We adopt the position that BHs are non-singular vacuum energy objects, cosmologically coupled, and produced solely by stellar collapse. To describe this process phenomenologically, we define the physical baryon density as follows:
$\rho_{b}:=\begin{cases}\displaystyle\frac{\rho_{\mathrm{cr}}\Omega_{b}}{a^{3}}&a<a_{i}\\ \displaystyle\frac{\rho_{\mathrm{cr}}\Omega_{b}}{a^{3}}-\frac{\Xi}{a^{3}}\int_{a_{i}}^{a}\psi\frac{\mathrm{d}a^{\prime}}{Ha^{\prime}}&a\geqslant a_{i}\end{cases}\,.$ (2.1)
Here, $\psi$ represents a model for the comoving star formation rate density (SFRD). The parameters $\rho_{\mathrm{cr}}$ and $\Omega_{b}$ denote the critical density and baryon fraction, respectively, inferred from early-universe measurements projected to the present day. In essence, the physical baryon density evolves as inferred by, for example, Planck, until the onset of stellar first light at $a_{i}$. After this point, a constant fraction $\Xi$ of baryons are depleted in synchrony with the assembly of stellar mass. The assumption that $\Xi$ is constant reflects the simplification that its time evolution is subdominant to the time evolution of $\psi$. We generically define the DE contribution through its pressure as:
$P_{\Lambda}:=w\rho_{\Lambda}\,,$ (2.2)
[…]
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}+\frac{3}{a}\rho_{\Lambda}(1+w)=\frac{\Xi}{Ha^{4}}\psi\,.$ (2.3)
In the limit where $\Xi\to 0$, DE evolves solely through gravitation
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}=-\frac{3}{a}\rho_{\Lambda}(1+w)\,.$ (2.4)
Given the $w_{0}w_{a}$ model in Eq. 1.1, this becomes
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}=-\frac{3}{a}\rho_{\Lambda}\left[(1+w_{0}+w_{a})-w_{a}a\right]\,.$ (2.5)
This equation is separable, and the solution is
$\rho_{\Lambda}=N\frac{e^{3aw_{a}}}{a^{3(1+w_{0}+w_{a})}}\,,$ (2.6)
where $N$ is a normalization constant. Note that, for $w_{0}+w_{a}<-1$, the solution can become “phantom,” growing from zero without any source. To reproduce the present-day DE density,
$N=\rho_{\mathrm{cr}}\Omega_{\Lambda}e^{-3w_{a}}\,,$ (2.7)
assuming Robertson-Walker (RW) flatness $\Omega_{\mathrm{K}}:=0$ to write
$\Omega_{\Lambda}=1-\Omega_{m}$ (2.8)
to a good approximation. The evolution of DE density under the $w_{0}w_{a}$ model, described by $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a$ in Eq. 2.5 can be expressed entirely by $a$ with Eq. 2.6. This derivative, for all DESI+CMB+SNIa most-likely $w_{0}$ and $w_{a}$ is shown in solid lines of Figure 1. The BH DE model is characterized by $w:=-1$ and $\Xi\neq 0$. In this setting, Eq. 2.3 becomes
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}=\frac{\Xi}{Ha^{4}}\psi\,.$ (2.9)
For simplicity, we approximate the Hubble expansion rate $H$ as follows: since $\psi\to 0$ as $z\to 0$, then $\mathrm{d}\rho_{\Lambda}/\mathrm{d}a\sim 0$ for $z\lesssim 1$. For $z\gtrsim 1$, the universe is matter-dominated and the impact of $\rho_{\Lambda}$ is small. Provided that baryon consumption is not too large, the typical expression for the Hubble rate
$H=H_{0}\sqrt{\Omega_{m}/a^{3}+(1-\Omega_{m})}\,,$ (2.10)
remains approximately valid. Substitution into Eq. 2.9 gives
$\frac{\mathrm{d}\rho_{\Lambda}}{\mathrm{d}a}=\frac{\Xi\psi(a)a^{-5/2}}{H_{0}\sqrt{\Omega_{m}+(1-\Omega_{m})a^{3}}}\,.$ (2.11)
This derivative is shown for all DESI+CMB+SNIa best-fit $\Omega_{m}$ and $H_{0}$ in the dotted lines of Figure 1. To ensure the correct DE density today is recovered
$\Xi=\rho_{\mathrm{cr}}\Omega_{\Lambda}\left[\int_{a_{i}}^{1}\frac{\psi(a)a^{-5/2}}{H_{0}\sqrt{\Omega_{m}+(1-\Omega_{m})a^{3}}}~{}\mathrm{d}a\right]^{-1}\,.$ (2.12)
The value of $\Xi$ for each DESI+CMB+SNIa measurement of $\Omega_{m}$ and $H_{0}$ is provided in the Figure 1 legends.
Figure 2: Impact of different SFRDs on the $w_{0}w_{a}$ model fit to the rate of production of dark energy. Shaded regions indicates the DESI+CMB+DESY5 $w_{0}w_{a}$ measurement with central value $w_{0}$. The $w_{a}$ values are taken uniformly within twice the $68\%$ confidence regions. The displayed $w_{0}w_{a}$ fits to the BH DE model with different SFRDs span a reasonable range in uncertainty. The Harikane SFRD [84] is a direct measurement, from rest-frame UV bright sources found by JWST, but may underestimate the total SFRD. The Trinca SFRD [85] is an estimate of the total SFRD from a semi-analytic model.
3 Results
We can now answer the question: If the production of DE is described by cosmologically coupled BH production, what conclusions can be drawn about DE evolution when the data are interpreted using a $w_{0}w_{a}$ model? To investigate this, we perform a least-squares fit of the $w_{0}w_{a}$ model to the BH DE model defined by Eqs. 2.11 and 2.12. We consider each DESI+CMB+SNIa measurement for $\Omega_{m}$ and $H_{0}$ and adopt the standard Madau and Fragos parametrization for $\psi$ [83]. The results, shown in Figure 1 with dashed lines, indicate a good match to the DESI measured solid lines across all three cases. The least-squares $w_{0}w_{a}$ fit to BH DE falls entirely within $1\sigma$ of the DESI most-likely $w_{0}$ and $w_{a}$ for all data sets. The normalization of the BH DE model is determined by $\Omega_{m}$ and $H_{0}$, yielding $\Xi\simeq 5$ in all three scenarios.
Next, we examine the sensitivity of our results to the SFRD. While the SFRD at $z\lesssim 4$ is relatively well-constrained, the SFRD at higher redshifts remains uncertain. In Figure 2, we compare the results using the Madau and Fragos SFRD model across all redshifts to other SFRD models informed from James Webb Space Telescope (JWST) results at high redshifts. The Harikane SFRD [84], based on UV-bright measurements, is expected to provide a lower-bound on SFRD at $z>5$ (see also [86]). The Trinca SFRD, a $z>4$ prediction from the semi-analytic Cosmic Archaeology Tool, is informed by recent JWST data and suggests a slower initial decline in SFRD as redshift increases[85, 87]. For this SFRD at $z\leqslant 4$, we use the Madau and Fragos SFRD scaled by 2$\sim$ 2$\times$ 2 × for continuity at the junction. For simplicity, we present only the DESI+CMB+DESY5 data set for comparison, adopting the central $w_{0}$ and twice the $68\%$ confidence region for $w_{a}$. The other SFRDs show consistency at $\sim 1\sigma$ and $\sim 2\sigma$ levels respectively, leading us to conclude that our results align with contemporary estimates of the high redshift SFRD. We note that DE production at $z>10$ is not well-parameterized by the $w_{0}w_{a}$ model, which must adopt a prior so that $\rho_{\Lambda}\to 0$ as $a\to 0$.
4 Discussion
Physically, $\Xi\sim 5$ implies that, on average, $5M_{\odot}$ of baryons convert into BH DE for every $1M_{\odot}$ of stellar material produced. From Eq. 2.1, the fraction of baryons remaining today, relative to stellar first light, is approximately:
$\rho_{b}(a=1)/\Omega_{b}\rho_{\mathrm{cr}}\sim 0.54\,.$ (4.1)
We have used $\Omega_{b}h^{2}=0.02205$ [88, Eqn. (24.6)], with $h^{2}=(H_{0}/100)^{2}$, where $H_{0}$ is measured in units of $\mathrm{km}~{}\mathrm{s}^{-1}~{}\mathrm{Mpc}^{-1}$ depends on the DESI dataset adopted. This result implies that approximately $\sim 45\%$ of baryons are missing at late times.
In fact, there has been a “missing baryons problem” for over a decade, with measurements indicating a shortfall of about $30\%$ [89]. Recent estimates pioneering fast-radio bursts (FRBs) report a value $\Omega_{b}=0.05^{+0.021}_{-0.025}$ at $95\%$ confidence level [90],555c.f. text below their Eqn. 8 in Methods, where their central value drops by $20\%$ at comparable uncertainties although direct constraints still indicate a shortfall of around $\sim 75\%$ [91, Fig. 1]. Our estimated survival fraction is marginally consistent with the FRB measurement, but we believe the actual consumption is likely lower. At face value, $\sim 45\%$ consumption would result in approximately a $\sim 10\%$ drop in $\Omega_{m}$ across $41<z<$ 4, enough to shift the onset of accelerated expansion earlier within DESI’s sensitivity. For consistency with DESI-1YR’s reported results, we do not incorporate this effect into Eq. 2.11, leaving it for future study.
This consumption of baryons may help explain other unexpected aspects of the DESI-1YR analysis. DESI-1YR results find a most probable summed neutrino mass $\sum m_{\nu}$ equal to zero [Fig. 11, Left Panel 8](see also [92]), contradicting the computed $\sum m>0.059~{}\mathrm{eV}$ based on $\Delta m_{ij}^{2}$ constraints from neutrino oscillation measurements [e.g. 93, 94]. [^6] This standard constraint follows from algebra $\sqrt{\Delta m_{21}^{2}}+\sqrt{\Delta m_{31}^{2}}=(m_{2}-m_{1})+(m_{3}-m_{1})<m_{2}+m_{3}<\sum m_{\nu},$ (4.2) under the assumption that $m_{\nu}\geqslant 0~{}\forall\nu$ and $m_{i}$ are monotonically increasing in $i$ (normal hierarchy). The summed neutrino mass approaching zero in DESI-1YR can be interpreted as a preference for lower matter density at late times, relative to that inferred from the CMB. [^7] It has also been proposed that a preferred zero summed neutrino mass could arise from new physics in the neutrino sector [95] We suspect that the baryon consumption required in the BH DE scenario will allow $\sum m_{\nu}$ to increase towards physically reasonable values. It is important to note that $\Omega_{b}h^{2}$ and $\sum m_{\nu}$ are not entirely degenerate, as the required DE density constrains the amount of baryon consumption.
Although the amount of baryonic density consumed is technically viable, the astrophysical processes directing it towards BHs are unclear. Assuming a Chabrier distribution of stellar masses at birth $\mathrm{d}N/\mathrm{d}m\propto m^{-2.3}$, the fraction of stellar mass in stars large enough to form BHs ($m\gtrsim 20M_{\odot}$) is $f=0.4$. If this fraction of stellar mass becomes BH mass, the value of $\Xi$ implies an additional growth factor of approximately 12.5$\sim$ 12.5$\times$ 12.5 × is required. We consider some channels that could contribute to this factor. The stellar Initial Mass Function (IMF) at $z\gtrsim 1$ is an active area of investigation [e.g. 96, 97]. If the IMF at $z\gtrsim 1$ were ‘top-heavy,’ as suggested by some studies [98, 99, 100, 101], the required factor decreases. For example, an IMF $\propto m^{-2.1}$ [e.g. 102] gives $f=0.74$, reducing the factor to 6.7×6.7$\times$6.7 ×. Post-collapse accretion could account for part of $\Xi$, though estimating its extent is challenging. Accretion onto isolated BHs is believed to be inefficient. Because $\Xi$ describes average behavior, a scenario where the high-end of the BH mass function accretes by a factor of $\sim 10^{2}\text{\textendash}10^{3}\times$, while the low end evolves passively, is conceivable. Gravitational processes can also produce mass. This physics is independent from $\Xi$ because Eq. 2.3 has no knowledge of local gravitational processes. The Press and Teukolsky process [103, 104] can convert BH spin into mass, provided radiated energy is absorbed by further infalling material during core collapse. Local strong gravitational effects would effectively amplify the term on the right-hand side of Eq. 2.3, without altering the depletion described by Eq. 2.1. We note that such effects may involve mediating fields, such as scalar fields [105, 106].
The “cosmic noon” scenario we consider here differs from the previously considered “cosmic dawn” scenarios, where high-$z$ BH production leveraged the $\sim 1/a^{3}$ factor in Eq. 2.9 to achieve $\Omega_{\Lambda}\rho_{\mathrm{cr}}$ from a small fraction of baryons [107, 61]. We note that any contribution from Population III stellar collapse and direct-collapse BHs at $10\lesssim z\lesssim 25$ would behave, at background order, as an effective cosmological constant. A scenario where $0.5\Omega_{\Lambda}$ is established at high-$z$ would, in principle, reduce $\Xi$ by an additional factor of two. Despite the agreement in shape, it is evident that further work is required to understand and estimate baryon consumption for a “cosmic noon” scenario.
In summary, DESI data suggest time-evolution in the dark energy density at $2.5\operatorname{-}3.9\sigma$ confidence when combining its Baryon Acoustic Oscillation measurements with Cosmic Microwave Background and various supernovae datasets. The time-evolution is characterized by a two-parameter dark energy equation of state $w=w_{0}+w_{a}(1-a)$, where $a$ is the scale factor. Our analysis shows that the DESI most-likely values for $w_{0}$ and $w_{a}$ are recovered when a $w_{0}w_{a}$ model is fit to dark energy produced by baryon conversion in cosmologically coupled black holes. Baryon conversion is constrained solely by measurements of the comoving star formation rate density, which are independent of the DESI measurements and previous observational studies of cosmologically coupled black hole solutions. As DESI gathers more data, the increased statistical precision will further constrain $w_{0}$ and $w_{a}$, potentially enabling DESI to measure the time evolution of dark energy independently, without relying on external Type Ia Supernova data.
Acknowledgments
KC thanks the Sakai Group at Yamaguchi University for their hospitality during the preparation of this work. KC further thanks Ryo Saito (Yamaguchi) for comments on $w_{0}w_{a}$ priors, Joel Weiner (Hawai‘i) for discussion, and Tom Browder (Hawai‘i) for comments on neutrinos. GT acknowledges support through DoE Award DE-SC009193. RAW acknowledges support from NASA JWST Interdisciplinary Scientist grants NAG5-12460, NNX14AN10G and 80NSSC18K0200 from GSFC. The work of NF is supported by DOE grant DOE-SC0010008. NF thanks the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452, and the Sloan Foundation for its partial support.